OpenStudy (anonymous):

anyone know how to help me? O_o f(1) = 2 and f(2) = 4, f(n) = f(1) + f(2) + f(n - 1), for n > 2 f(5) = ______

OpenStudy (anonymous):

Hello @PIRATE..LORD,This is INeedHelpPlease assisting you today! You have to find f(5) And given that, \[f(n) = f(1) + f(2) + f(n - 1)\] Put n=5 \[f(5)=f(1)+f(2)+f(5-1)\] f(1) = 2 and f(2) = 4, Substitute the values, \(\LARGE f(5)=2+4+f(4)\)-------Let it be equation number 1 Now we need f(4) Again use, \[f(n) = f(1) + f(2) + f(n - 1)\] f(4)=f(1)+f(2)+f(3) \(\LARGE f(4)=2+4+f(3)\)-----------Let it be equation number 2 Now we need f(3) Again and again use, \[f(n) = f(1) + f(2) + f(n - 1)\] \[\LARGE f(3)=f(1)+f(2)+f(2)\] f(3)=2+4+4 f(3)=10 Now that we have f(3)'s value,put it in equation number 2 It will reduce to,f(4)=6+f(3)=6+10=16 Now that we have f(4)'s value,put it in equation number 1 It will reduce to, \[\Huge f(5)=6+f(4)=6+16=22\]. That is your final answer!